OpenStudy (saifoo.khan):

Domain & Range of: -0.035 x^2+1.4 x+6.

OpenStudy (turingtest):

the legend asks a question!

OpenStudy (saifoo.khan):

And Superstar answers it. ;)

OpenStudy (turingtest):

i didn't say I could but i'll try

OpenStudy (turingtest):

domain is all reals clearly

OpenStudy (saifoo.khan):

range?

OpenStudy (turingtest):

derivatives to find max and mins you think?

OpenStudy (saifoo.khan):

No idea. :/

OpenStudy (anonymous):

the domain is all real x: and the curve is concave so find the maximum value. so range is -infinity, maximum

OpenStudy (saifoo.khan):

what do u mean dipan?

OpenStudy (saifoo.khan):

Elaborate.

OpenStudy (turingtest):

after derivative -.07x+1.4=0 x=-20 plug in to original equation...

OpenStudy (saifoo.khan):

this is another form of the same equation: -(0.035) (x - 20)² + 20

OpenStudy (saifoo.khan):

Ok, solve this to completing square form, SHOW STEPS as well: -0.035 x^2+1.4 x+6.

OpenStudy (saifoo.khan):

-35x^2 + 1400 + 6000

OpenStudy (anonymous):

ALL REAL NUMBERS

OpenStudy (saifoo.khan):

what?

OpenStudy (turingtest):

-0.035 x^2+1.4 x+6=f(x) f''(x)=-.07 (concave down) f(-20)=-36 is the max, so the range is (-infinty,-36)

OpenStudy (anonymous):

f(x)=-0.035 x^2+1.4 x+6

OpenStudy (saifoo.khan):